OpenStudy (anonymous):

Perfect Square Formula; (v+8)^2

OpenStudy (angela210793):

\[(a+b)^{2}=a ^{2}+2ab+b ^{2}\]

OpenStudy (angela210793):

can u try this urself???

OpenStudy (anonymous):

hold up

OpenStudy (angela210793):

take ur time :)

OpenStudy (anonymous):

(v+8)^2 = v^2 + 2 x V x 8 + 8^2

OpenStudy (anonymous):

but then i don't know how to simplify it:/

OpenStudy (angela210793):

yes, good job :) =v^2+16v+64 :)))

OpenStudy (anonymous):

thank you!

OpenStudy (angela210793):

ur welcome :)

OpenStudy (anonymous):

could you do another one with me? and like check it

OpenStudy (angela210793):

sorry...wasn't here...but yea..I'd love to help you :)